OpenStudy (tester97):

Someone help???): Suppose a parabola has an axis of symmetry at x = –7 , a maximum height of 4 and also passes through the point (–6, 0). Write the equation of the parabola in vertex form. Show all work.

OpenStudy (shamil98):

Do you know the vertex form?

OpenStudy (anonymous):

0 = a(-6 + 7)^2 + 4 a = -4 y = -4(x+7)^2 + 4

OpenStudy (tester97):

@xxferrocixx how did you get that answer? could youu break it down a little bit more?

OpenStudy (shamil98):

The maximum height is at 4, and the axis of symetery is at x= -7, vertex form is y = a(x-h)^2 + k \[(h,k) -> (-7,4)\] Passes at (0,-6) 0 = a(-6+7)^2 + 4 = a => -4 so the parabola is \[y = -4(x+7)^2 + 4\]

OpenStudy (tester97):

thanks shammy :)